John has $150 in his piggy bank and plans to add $2 each week. Sabrina has $200 in her piggy bank and plans to add $6 each week. After how many weeks will Sabrina’s piggy bank have twice as much money as John’s piggy bank.
Answer by addingup(3677)

How many weeks? We don't know. Let's call it x
2(150 + 2x) = 200 + 6x
300 + 4x = 200 + 6x
100 = 2x
x = 100/2 = 50
At 50 weeks Sabrina’s piggy bank will have twice as much money as John’s piggy bank
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
2(150 + 2(50)) = 200 + 6(50)
300 + 200 = 200 + 300 Add each side and you'll see the answer is correct.
.
Happy learning